# Leveraging Domain Knowledge for Efficient Reward Modelling in RLHF: A Case-Study in E-Commerce Opinion Summarization

### v1

@@ Appendix F Statistics of the OpinPref dataset
Table 8: Statistics of the OpinPref dataset. We use NLTK to split into tokens.
Appendix G Annotator Details
@@
We include two disjoint sets of annotators in our work – first for creation of OpinPref ($3$ annotators), second for human evaluation  ($3$ annotators). For both annotations, we use domain experts. The domain experts are NLP researchers (age group: $24-30$) who have worked in Opinion Summarization for a long time, with publication experience (in $*{}^{*}$ conferences). The domain experts for human evaluation also have a similar profile. The annotators have been paid generously, based on the standard annotation rates in the geographical location.
Appendix H All Evaluation Results
@@
We include all evaluation results in this section. In Tables 9, 10, 11 and 11 we include pairwise comparison results, in a win/tie/loss format. We also include results on evaluation on the features in Tables 13, 14 and 15.
@@
Inductive-Bias $\mathbf{0.56}/0.09/0.28$ $\mathbf{0.46}/0.18/0.31$ $0.56/0.12/0.28$
@@
Table 9: Pairwise Win/Tie/Loss Results for all models in Human Evaluation for Amazon Benchmark. We format the data as: win/tie/loss.
@@
Table 10: Pairwise Win/Tie/Loss Results for all models in Gpt-$4$ Evaluation for Amazon Benchmark. We format the data as: win/tie/loss.
@@
Table 11: Pairwise Win/Tie/Loss Results for all models in Gpt-$4$ Evaluation for Flipkart Benchmark. We format the data as: win/tie/loss.
@@
Table 13: Intrisic Evaluation results on the Amazon Benchmark for all the models. Legend: AC: aspect-coverage, OPF: opinion-faithfulness, OPC: opinion-coverage, CC: conciseness, RL: relevance, HL: hallucination, LC: language-correctness.
@@
Table 14: Intrisic Evaluation results on the Flipkart Benchmark for all the models. Legend: AC: aspect-coverage, OPF: opinion-faithfulness, OPC: opinion-coverage, CC: conciseness, RL: relevance, HL: hallucination, LC: language-correctness.
@@
Table 15: Intrisic Evaluation results on the Oposum+ Benchmark for all the models. Legend: AC: aspect-coverage, OPF: opinion-faithfulness, OPC: opinion-coverage, CC: conciseness, RL: relevance, HL: hallucination, LC: language-correctness.

### v2

@@ Appendix H Statistics of the OpinPref dataset
Table 12: Statistics of the OpinPref dataset. We use Nltk to tokenize the text.
Split Characteristic $\mu$ $\sigma$
train # reviews $13.24$ $10.07$
# summaries $8.90$ $0.34$
# words in review $49.0$ $10.78$
# words in summary $78.28$ $34.45$
validation # reviews $10.53$ $6.80$
# summaries $8.98$ $0.16$
# words in review $48.65$ $10.63$
# words in summary $74.26$ $34.27$
Table 13: Statistics of PromptOpinSumm dataset. We use Nltk to tokenize the text.
Reviews: i really like these boots. they can be a beast to get on, like any boot fit to last; once on, they are incredibly comfortable. I have had them for 4 years and they still look great - worn in but not worn out… What great boots! They do take some breaking in - like all real leather does. Run maybe a tad small in the general spectrum, but hey, remember the break in period. Look fabulous, will last me for the next twenty years…. They bridge the fashion vs. function quandary beautifully. Every woman needs a pair of these Frye boots! Fit is true to normal size. I wear a standard 9. These fit in the shaft for any ladies who have bigger calves as I do. You have to fight to get them on first few times and they wear stiff until fully broken in. Worth the money and break in time. Boots excellent quality, beautiful skin, a modern model. But! they are very narrow! This is the second pair of shoes is a brand and my 8.5 size fits me perfectly, even he little more than usual. But these are very narrow! go to a cobbler, let him try to stretch. Be careful! bought these for the wife to have something to kick around in and keep up with the kids, but she doesn’t like them. Her heel lifts with every step, and they’re stiff and heavy. Not an everyday casual shoe. Don’t get me wrong, they’re really good boots, but they used to make them better. My last pair from a few years ago were much higher quality. Thicker sole, better full grain leather, now they only make the black in ’grease’, which stains all my socks black. Keeping them, but still disappointed in Frye. i bought it just for the heavy snow cuz my UGG got wet through.even though it is a little expensive, but the boots are nice and it was shipped to me for only two days.!!!!!! i love it They were comfortable the moment I put them on. The sizes are exact. The color the way they are made are what I expect from Frye which I’ve been wearing since the 70’s. With a 20year break from not wearing Frye it was a pleasant surprise the quality has stood the test of time.
Supervised: These Frye boots are a popular choice for women due to their comfort and durability. However, they can be a bit challenging to put on and require a break-in period. Some users have reported that they run small in size and are narrow, requiring a cobbler’s assistance for a proper fit. The leather quality is generally good, but some users have noted that the leather may not last as long as they had hoped.
NaiveMean: Users generally appreciate the comfort and durability of the Frye boots, with some mentioning the need for a break-in period. Sizes run true to normal, but some users find them to be a bit narrow. The boots are praised for their stylish appearance and ability to bridge the fashion and function quandary. However, some users have noted that they take some time to break in and are not suitable for everyday casual wear.
Synth-Feedback: These Frye boots are praised for their comfort and durability, with some users noting the need for a break-in period. However, some users have noted that they run small in size and require stretching, while others find them narrow and stiff. The leather is described as beautiful and of good quality, with a modern model. Some users have had them for several years and have not worn them out.
Inductive-Bias: These Frye boots are praised for their comfort and durability, with users expressing satisfaction with their fit and appearance. However, some users have noted that they require a break-in period and can be a bit narrow. The leather quality is praised for its durability and modern model, with some users expressing disappointment with the lack of improved quality in recent years.
Table 14: Example generation (randomly sampled) for some input reviews from all the models. Olive implies faithful/correct generation, while red indicates hallucinated text, or repetition. We see that only Inductive-Bias is free from red text. The model closest in performance to Inductive-Bias, the NaiveMean model, misses out on two aspects: leather-quality and quality-degradation. Inductive-Bias covers both, while being concise. We do not include DPO model in this comparison, as it was too verbose.
Appendix I Annotator Details
@@
We include two disjoint sets of annotators in our work—first for creation of OpinPref ($3$ annotators), second for human evaluation  ($3$ annotators). For both annotations, we use domain experts. The domain experts are NLP researchers (age group: $24-30$) who have worked in Opinion Summarization for a long time, with publication experience (in A/A conferences). The domain experts for human evaluation also have a similar profile. The annotators have been paid generously, based on the standard annotation rates in the geographical location.
Appendix J All Evaluation Results
@@
We include all of the evaluation results in this section. In Tables 15, 16, 17 and 17 we include pairwise comparison results, in a win/tie/loss format. We also include results on evaluation on how the models perform on the domain features in Tables 19, 20 and 21.
@@
Inductive-Bias $\mathbf{0.56}/0.09/0.28$ $\mathbf{0.46}/0.18/0.31$ $\mathbf{0.56}/0.12/0.28$
@@
Table 15: Pairwise Win/Tie/Loss Results for all models in Human Evaluation for Amazon benchmark. We format the data as: win/tie/loss, win specifies how many time the row won over the column.
@@
Table 16: Pairwise Win/Tie/Loss Results for all models in Gpt-$4$ Evaluation for Amazon benchmark. We format the data as: win/tie/loss, win specifies how many time the row won over the column.
@@
Table 17: Pairwise Win/Tie/Loss Results for all models in Gpt-$4$ Evaluation for Flipkart benchmark. We format the data as: win/tie/loss, win specifies how many time the row won over the column.
@@
Table 19: Intrisic Evaluation results on the Amazon benchmark for all the models. Legend: AC: aspect-coverage, OPF: opinion-faithfulness, OPC: opinion-coverage, CC: conciseness, RE: relevance, HL: hallucination, LC: language-correctness.
@@
Table 20: Intrisic Evaluation results on the Flipkart benchmark for all the models. Legend: AC: aspect-coverage, OPF: opinion-faithfulness, OPC: opinion-coverage, CC: conciseness, RE: relevance, HL: hallucination, LC: language-correctness.
@@
Table 21: Intrisic Evaluation results on the Oposum+ benchmark for all the models. Legend: AC: aspect-coverage, OPF: opinion-faithfulness, OPC: opinion-coverage, CC: conciseness, RE: relevance, HL: hallucination, LC: language-correctness.